OpenStudy (anonymous):

Help integrate: cot^3(x)

OpenStudy (anonymous):

table that one

OpenStudy (anonymous):

∫ tan³(x) dx = ∫ tan(x)tan²(x) dx = ∫ tan(x)(sec²(x) - 1) dx = ∫ tan(x)sec²(x) - tan(x) dx = ∫ tan(x)sec²(x) dx - ∫tan(x) dx = ∫ tan(x)sec²(x) dx - ln|sec(x)| let u = tan(x) du = sec²(x) dx = ∫ u du - ln|sec(x)| = u²/2 - ln|sec(x)| + C = tan²(x)/2 - ln|sec(x)| + C

myininaya (myininaya):

\[\text{ recall } \cot^2(x)+1=\csc^2(x)\]

myininaya (myininaya):

=>\[\cot^2(x)=\csc^2(x)-1\] now follow what indy did above it will the same process for this integral

OpenStudy (anonymous):

how did you cange it to tan^3 x

myininaya (myininaya):

he didn't he was giving you an example to follow

myininaya (myininaya):

\[\int\limits_{}^{}\cot(x)\cot^2(x) dx=\int\limits_{}^{}\cot(x)(\csc^2(x)-1) dx\]

myininaya (myininaya):

\[=\int\limits_{}^{}\cot(x)\csc^2(x) dx-\int\limits_{}^{}\cot(x) dx\]

OpenStudy (anonymous):

okay, i got that far, myininaya, but waht do i do next

myininaya (myininaya):

for the first one let u=cot(x)=>du=-csc^2(x) dx second one let u=sin(x)=> du=cos(x) dx

myininaya (myininaya):

and i like u you can call your substitution something different for the second one if you like

OpenStudy (anonymous):

okay, that actually now is starting to sound Good, let me try it and i will get back to you

OpenStudy (anonymous):

i also like u

myininaya (myininaya):

i mean i don't like u lol i mean i don't know you i mean u as in the substitution u lol

OpenStudy (anonymous):

as in u the variable

OpenStudy (anonymous):

you dont like me :(

OpenStudy (anonymous):

I like u too

OpenStudy (anonymous):

hilarious one @ mathsnake

myininaya (myininaya):

lol stupid u not you

myininaya (myininaya):

i don't like u anymore

OpenStudy (anonymous):

lol dummy U

myininaya (myininaya):

its confusing

OpenStudy (anonymous):

as in dummy variable

myininaya (myininaya):

right